### v1

# R+X: Retrieval and Execution from Everyday Human Videos

Abstract
We present R+X, a framework which enables robots to learn skills from long, unlabelled, first-person videos of humans performing everyday tasks. Given a language command from a human, R+X first retrieves short video clips containing relevant behaviour, and then executes the skill by conditioning an in-context imitation learning method on this behaviour. By leveraging a Vision Language Model (VLM) for retrieval, R+X does not require any manual annotation of the videos, and by leveraging in-context learning for execution, robots can perform commanded skills immediately, without requiring a period of training on the retrieved videos. Experiments studying a range of everyday household tasks show that R+X succeeds at translating unlabelled human videos into robust robot skills, and that R+X outperforms several recent alternative methods. Videos are available at https://www.robot-learning.uk/r-plus-x.
**footnotetext: Equal contribution, random order.**footnotetext: Contact at: {g.papagiannis21, n.di-palo20}@imperial.ac.uk
Figure 1: Given a language prompt, R+X first retrieves short relevant video clips extracted from a long unlabelled video of a human performing everyday tasks, recorded with a wearable camera. With the use of the retrieved video clips and a VLM, R+X performs in-context imitation learning allowing it to immediately generate and execute the desired behaviour on the robot.
1 Introduction
Robot learning of diverse, everyday tasks in natural environments, is a significant challenge. Imitation learning is a promising solution which has been widely adopted in recent years [1, 2, 3], but it remains difficult to scale up due to the cost of hardware and human demonstration time. If robots could instead learn from data which does not require any robot hardware or dedicated human time, we may be able to scale up data collection and learning substantially.
In this work, we study the problem of learning robot skills from long, unlabelled, first-person videos of humans performing everyday tasks in everyday environments. "Long, unlabelled videos" means that a human simply goes about their everyday life and passively records videos of themselves, without the need to specify which behaviour is being performed. "Everyday environments" means that videos contain diversity in tasks, scenarios, objects, distractors, illumination, and camera viewpoints and motion. It is likely that such videos will become abundant through the adoption of wearable devices such as AR headsets and glasses [4, 5, 6, 7], and thus this offers a significant opportunity for future scalability if robots could learn from such videos.
Previous approaches to learning from videos of humans have often relied on a set of strong constraints and assumptions, such as human videos manually aligned with robot videos, human videos manually labelled with language descriptions, or demonstrations on robot or mocap hardware [8, 9, 1, 10, 11, 12, 13]. In this work, we remove all of these constraints, and present a framework that requires only an unlabelled first-person video of a human performing tasks.
Our framework, R+X, is a two-stage pipeline of Retrieval and EXecution (see Figure 1), which uses Foundation Models for both stages. Upon receiving a language command from a human, a Vision Language Model retrieves all clips where the human executes the specified task. By extracting the trajectories of the human’s hand in each clip, we then employ a few-shot in-context imitation learning method to condition on these trajectories, which enables a robot to ingest, learn from, and replicate the retrieved behaviours, whilst adapting to previously unseen settings and objects. Recent literature has demonstrated that, by finetuning large VLM on robotics data, they can transfer their common knowledge and ability to reason and plan to robotics settings [14, 15]. This, however, requires very expensive finetuning of often intractably large models. With our proposed framework, we can equally leverage these abilities, but now via retrieval and video understanding, thus without the need for any finetuning. Rather than requiring explicit policy training, the retrieved skills are executed immediately following the language command. We later demonstrate the benefits that this brings over spatial and language generalisation, compared to large monolithic policy networks.
In summary, the two main properties of R+X are: 1) it enables robots to execute everyday tasks from language commands, given long, unlabelled, first-person videos, collected naturally by recording a human’s daily activities; and 2) it achieves this without the need for any training or finetuning of models, allowing a robot to learn and execute tasks immediately after the language command.
2 Related Work
Figure 2: The main assumptions and constraints of many recent Learning from Observation methods.
Collecting Robotics Data. Scaling data collection has proven to be a successful path towards increasingly more general machine learning models [16, 17, 18, 19]. To collect robotics data, a popular paradigm is through kinesthetic teaching or teleoperation, collecting datasets of paired observations and actions [3]. This, however, needs dedicated robot hardware and optionally teleoperation hardware, and needs human operators to allocate their time to proactively teach a robot new tasks. In our framework, humans interact with their environments much more naturally, collecting data passively during their daily activity, resulting in a more scalable and time efficient paradigm.
Learning from Human Videos. Many recent works have proposed solutions to teach new skills to robots by observing a human executing such tasks, known as Learning from Observation. However, as we illustrate in Fig. 2, they often rely on a set of assumptions and constraints that limit their use "in-the-wild". Bahl et al. [9], Wang et al. [20] require a combination of human data and either robot exploration or teleoperation, therefore needing active assistance by humans in teaching the robot, and videos are recorded from a fixed, third-person camera. R+X relies entirely on the videos recorded by the human with a mobile camera in their natural environments. Heppert et al. [10], Zhu et al. [12], Bahety et al. [11] can learn robot skills entirely from human videos. Heppert et al. [10], Zhu et al. [12], however, they focus on replicating the object trajectory from the demonstration, and therefore cannot perform non prehensile tasks that do not involve grasping and moving, such as pushing or pressing, which R+X can execute. Bahety et al. [11] can learn a large repertoire of skills, but still relies on learning single tasks in isolation from a fixed camera, while R+X can replicate tasks from a given language command after receiving a single, long, unlabelled video depicting tens of tasks, without the need to specify which clip demonstrated which behaviour. Methods like Wang et al. [13] allow a human to naturally interact with their environment while collecting data that can teach a robot new skills. However, they require additional hardware to wear, like specialised MoCap gloves. R+X only needs a single RGB-D camera, and does not require any extrinsics calibration. This means that data could be recorded using a wearable camera, smart glasses, AR visors, and more.
Language-Conditioned Policy Learning. Robots should learn a wide repertoire of skills, and be able to execute many different tasks when deployed. Language is considered a viable way to instruct robots and guide their task execution [15, 14, 21]. The common approach is to train a large, language-conditioned and image-conditioned policy that can, at test time, receive human commands and camera observations [3]. This, however, presents some problems: training such models can require enormous computational effort [14], and yet robots should ideally continually learn new skills over time, where re-training or finetuning such networks becomes problematic. R+X does not train or finetune any networks, but takes advantage of pre-trained models able to retrieve examples of relevant tasks [17], enabling robots to learn and execute new behaviour directly at deployment via in-context learning [22].
Figure 3: Upon receiving a language command, R+X retrieves all the relevant clips from the human video. Each retrieved clip is transformed from pixels to a sparse 3D points representations of the hand joints movement and salient parts of the visual observation.
3 R+X: Retrieval and Execution
Figure 4: The visual 3D keypoints of the first frame of each of the Z videos obtained from the retrieval phase along with each extracted hand joint trajectory are used as context for KAT [22], a recent in-context imitation learning method. To execute a skill, visual 3D keypoints are extracted from the live observation and used as input to an imitation learning model which generates a sequence of hand joints. By mapping this sequence to gripper poses the robot executes the desired task.
We now describe R+X, our proposed method to learn robot skills from long, unlabelled videos of humans interacting with their environments. We assume R+X has access to a long, unlabelled video of a human performing a multitude of tasks in many different locations. We call this long, first-person video of everyday tasks the "human video" $\mathcal{H}$ in the rest of the paper. The goal of our method is to learn to emulate the behaviours recorded in the human video upon receiving a language command from the human.
Overall, from a high-level perspective, our pipeline takes three inputs during deployment: 1) a single, long, unlabelled video of all the interactions recorded by the humans, $\mathcal{H}$; 2) a task to execute in language form, $\mathcal{L}$; and 3) the robot’s current observation as an RGB-D image, $\mathcal{O}_{live}$. It then outputs a trajectory of 6-DoF gripper poses, which are executed by the robot to perform the commanded task.
There are however a set of non-trivial challenges. While the robot receives a language command at deployment, it receives no language information before deployment: it only has access to the unlabelled video. Additionally, the robot receives no action information: unlike the case of teleoperation, no joints or end-effector position/velocity/acceleration data are recorded, and the correct movements need to be inferred by the videos alone; this problem is additionally complex due to the cross-embodiment between human videos and robot actions. Furthermore, as the human is interacting with their natural environment, the visual observations can be filled with distractor objects typical of everyday environments, which are unrelated to the task the human is performing. To tackle all these aforementioned challenges, we leverage the abilities of Foundation Models.
Specifically, at the first phase of R+X, which we call the retrieval phase, we use a VLM to extract from the human video all the examples of the desired behaviour described in the prompt $\mathcal{L}$, as a list of $Z$ shorter video clips $[\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{Z}]$. We map the $Z$ videos into a lower dimensional 3D representation, extracting for each video: (1) a list of $K$ visual 3D keypoints that describe the scene, $[k_{1},k_{2},\dots,k_{K}]$ where $k=(x_{k},y_{k},z_{k})$, and (2) the movement of the human’s hand as a trajectory of length $T$ of 21 hand joints that parametrise the MANO hand model [23], $\mathcal{J}=[j_{1},j_{2},\dots,j_{T}]$, where $j=[[x_{0},y_{0},z_{0}],\dots,[x_{21},y_{21},z_{21}]]$. Finally, at the second stage of R+X, which we call the execution phase, to emulate the behaviours observed in the retrieved video clips, we condition a few-shot in-context imitation learning model [22] on this data which, given a live observation of the environment, generates a trajectory of 3D hand joints to execute the desired behaviour described in the prompt. To map such joints to gripper poses, we designed a heuristic that we describe in the Supplementary Material.
3.1 Retrieval: Extracting visual examples from a long, unlabelled video
The first main phase of R+X is retrieval, shown in Figure 3. Upon receiving a language command $\mathcal{L}$, and given the human video $\mathcal{H}$, the goal is to retrieve all video clips from the human video that depict the execution of the requested task. This is accomplished using a recent VLM, Gemini Pro 1.5 Flash [24]. Gemini, which we denote $\mathcal{G}$, is natively multi-modal and can take as input images, videos, and text, and outputs text. We prompt the model with the human video, and ask it to retrieve the starting and ending seconds at which the received task happens. The inputs of this phase are therefore a language command and the human video, and the output is a list of $Z$ shorter video clips demonstrating the desired task, $\mathcal{G}(\mathcal{H},\mathcal{L})\rightarrow[\mathcal{V}_{1},\dots,\mathcal{V}_{Z}]$. Each clip comprises $T$ RGB-D frames, where $T$ can vary across different clips.
3.1.1 Preprocessing videos into a sparse 3D representation
Given the $Z$ extracted video clips, $[\mathcal{V}_{1},\dots,\mathcal{V}_{Z}]$, we apply a preprocessing step that converts each video clip from a list of RGB-D frames to a set of 3D points describing the visual scene and the trajectory of the human’s hand, a representation that we will then feed to our few-shot in-context imitation learning model.
Visual 3D keypoints. To transform the complex RGB-D observations into a lower-dimensional, easier to interpret input, we harness the powerful vision representations generated by DINO [25], a recent Vision Foundation Model. As proposed in [26, 22], given the $Z$ clips retrieved as described before, we find a set of $K$ common 3D visual keypoints that capture interesting semantic or geometrical aspects of the scene. We extract these keypoints from the first frame of each of the $Z$ videos only.
Figure 5: Visual keypoints are extracted autonomously from RGB via a matching algorithm [26].
We first compute the DINO descriptors for the first frame of each of the $Z$ videos, obtaining a list of $Z$ different $N\times 384$ outputs, where $N$ is the number of patches of each image [27, 28, 25]. Via a clustering and matching algorithm [26], we select from the $N\times 384$ descriptors of the first frame of the first video, $\mathcal{O}_{\mathcal{V}_{1},1}$, a list of the $K$ descriptors that are the most common in all the remaining $Z-1$ frames, as shown in Figure 5. We denote these descriptors as $\mathcal{D}\in\mathbb{R}^{K\times 384}$. Therefore, this way, we autonomously extract descriptors that focus on the object of interest, as its appearance is common among videos, while distractors and overall scene will vary [26, 22].
Finally, for each of the $K$ descriptors in $\mathcal{D}$, we extract keypoints by finding the $K$ nearest neighbours between the $N\times 384$ descriptors of each of the $Z$ frames, and compute their 2D coordinates. We then project these in 3D using each frame’s depth image and known camera intrinsics. As such, given the retrieved clips, we obtain and store a list $\Lambda=[\mathcal{K}_{V_{1}},\dots,\mathcal{K}_{V_{z}}]\in\mathbb{R}^{Z\times K\times 3}$ of visual 3D keypoints, $K$ for each clip, where each $\mathcal{K}_{V_{i}}=[k_{1},...,k_{K}]$ and $k_{j}=(x_{j},y_{j},z_{j})$. For a more detailed description please refer to Amir et al. [26].
Hand Joint Actions. To extract human actions from each retrieved video clip we use the HaMeR model [29], a recent technique for 3D hand pose estimation from images based on DINO. In particular, using HaMeR we extract the 3D hand pose from each video frame $\mathcal{O}_{\mathcal{V}_{z},t}$ (where $1\leq t\leq T$) of each of the $Z$ clips, represented as a set of 21 3D points, $j_{\mathcal{V}_{z},t}$, describing the hand joints that parameterise the MANO hand model, as it is commonly done in the literature [23]. As the camera moves between frames, we design a stabilisation technique to compute transformations between camera poses that is robust to dynamic scenes (for more details please see our Supplementary Material). This enables us to express the extracted hand joints relative to a single reference frame, that of the first camera frame of each clip. For each video clip $\mathcal{V}_{z}$, this process results in a sequence of 3D hand joint actions $\mathcal{J}_{\mathcal{V}_{z}}=[j_{\mathcal{V}_{z},1},\dots,j_{\mathcal{V}_{z},T}]$, expressed relative to the first frame of each video $\mathcal{O}_{\mathcal{V}_{z},1}$ where the visual 3D keypoints are also expressed in. As such, at the end of this process we are left with a list $\mathcal{M}=[\mathcal{J}_{\mathcal{V}_{1}},...\mathcal{J}_{\mathcal{V}_{Z}}]$ of $Z$ hand joint action sequences.
In summary, from the $Z$ retrieved video clips, we extract the list of hand joints actions $\mathcal{M}$, along with the list of visual 3D keypoints $\Lambda$, that will be used as context for our in-context imitation learning model, as $Z$ input-output pairs.
3.2 Execution: Few-Shot, In-Context Imitation from Video Examples
The second main phase of R+X is execution. Our framework is based on the use of a model capable of performing few-shot, in-context imitation learning, receiving a few examples of desired inputs and outputs pairs describing a desired behaviour, and able to replicate such behaviour immediately upon receiving a new input. We use Keypoint Action Tokens (KAT) [22] to achieve this, a recently proposed technique that takes 3D visual keypoints as input and outputs a trajectory of 3D points describing the gripper movement. Instead of explicitly training a model on robot data, KAT demonstrates that recent, off-the-shelf Large Language Models are able to extract such numerical patterns, and behave as few-shot, in-context imitation learning machines, without the need for any further finetuning.
Given the output of the retrieval phase $[\Lambda,\mathcal{M}]$ and a new, live RGB-D observation collected by the robot $\mathcal{O}_{live}$, we extract its visual 3D keypoints representation $\mathcal{K}_{live}$ by first extracting its $N\times 384$ DINO descriptors, and then finding the $K$ nearest neighbours to each of the $K$ descriptors in $\mathcal{D}$, that we obtained in the retrieval phase. This results in $K$ 2D coordinates, that we project in 3D. We then input to KAT as context $[\Lambda,\mathcal{M}]$, as $Z$ examples of the desired input-output mapping, and the new visual 3D keypoints, and generate a new trajectory of desired hand joint actions, $\text{KAT}([\Lambda,\mathcal{M}],\mathcal{K}_{live})\rightarrow\mathcal{J}_{live}$ with $\mathcal{J}_{live}=[j_{1},\dots,j_{T}]$, as shown in Figure 4. We then map the predicted trajectories of hand joints actions into gripper poses as described in our Supplementary Material, and execute them.
To summarize, given a language command $\mathcal{L}$ and a human video $\mathcal{H}$, R+X first retrieves $Z$ videos depicting the described behaviour using a VLM. From the retrieved videos, a list of $[\Lambda,\mathcal{M}]$ visual 3D points and hand joint actions are extracted. Then, to execute the desired behaviour described in the prompt $\mathcal{L}$, $[\Lambda,\mathcal{M}]$ along with the visual 3D keypoints of the live observation $\mathcal{K}_{live}$ are used as context for KAT, which performs few-shot in-context imitation learning to generate a trajectory of hand joints actions, which is then mapped to gripper poses and executed on the robot.
Figure 6: We test R+X on 12 everyday tasks, executed by a human in different rooms and with different distractors.
4 Experiments
Human Video. We collect the human video $\mathcal{H}$ using an Intel RealSense 455, worn by the human on their chest as shown in Figure 1. To reduce downstream computational time, we filter out each frame in which human hands are not visible right after recording. As our robot is single-armed, we limit ourselves to single-hand tasks. However, our method could be applied to bimanual settings and dexterous manipulators.
Robot Setup. At execution, we use a Sawyer robot equipped with a RealSense 415 head-camera and a two-fingered parallel gripper, the Robotiq 2F-85. As the robot is not mobile, we set up different scenes in front of it with variations of the tasks recorded by the human, placing several different distractors for each task, while the human video was recorded in many different rooms. Although we have a wrist-camera mounted to the robot as shown in some of our figures, we do not use it in our work.
Tasks. To evaluate our proposed framework, we use a set of 12 everyday tasks, where the human interacts with a series of common household objects, listed in Fig. 6. We include movements like grasping, opening, inserting, pushing, pressing, and wiping.
Baselines. We compare R+X, and its retrieval and execution design, to training a single, language-conditioned policy. To obtain language captions from the human video, we use Gemini to autonomously caption snippets of the video, obtaining a (observation, actions, language) dataset. We finetune R3M (ResNet-50 version [30]) [31] and Octo [32] on this data. We extend R3M to also encode language via SentenceBERT and use a Diffusion Policy [33] head to predict actions from intermediate representations. We denote this version as R3M-DiffLang.
More details on all aspects are provided in the Supplementary Material.
Method / Task Plate Push Wipe Beer Wash Box Kettle Micro. Basket Phone Can Light Avg.
R3M-DiffLang 0.5 0.7 0.4 0.7 0.5 0.5 0.4 0.8 0.7 0.4 0.7 0.3 0.55
Octo 0.5 0.8 0.5 0.6 0.5 0.5 0.4 0.7 0.6 0.4 0.6 0.3 0.53
R+X 0.6 0.8 0.7 0.8 0.6 0.7 0.6 0.8 0.7 0.7 0.8 0.6 0.7
Table 1: Success rates of the various methods on the 12 proposed tasks.
4.1 Results
Can R+X learn robot skills from long, unlabelled videos? How does it perform with respect to a monolithic language-conditioned policy? In these experiments, we evaluate the performance of R+X in learning a set of everyday skills. At deployment, we place the object to be interacted with in front of the robot, and issue a language command. We then run the retrieval and execution pipeline using the human video, the current observation and the language command. We run 10 episodes per task, for each episode randomising 1) the object pose, 2) the type and number of distractors, 3) for tasks where it is possible, we swap the object to interact with, with another one from the same class to test for generalisation (e.g. a different can, a different piece of clothing, a different telephone). More details are provided in the Supplementary Material.
Figure 7: Examples of hard spatial and language generalisation. Gripper moves from red to blue.
In Table 1, we report the performance of R+X on these tasks, together with the baselines. As we demonstrate, the framework is able to tackle a wide range of everyday tasks, surpassing the monolithic policy baselines. These results prove the benefit of modeling language-conditioned learning from observation as distinct retrieval and execution phases, to fully leverage the abilities of recent Foundation Models.
What are the main sources of difference in performance between R+X and a monolithic policy? In these experiments, we investigate more in detail what changes in the inputs lead to the most noticeable difference in performance between R+X and the baselines, R3M-DiffLang and Octo. We explored two aspects, related to two properties of R+X:
Hard Spatial Generalisation: R+X, by retrieving videos of the desired task, can also extract a series of relevant keypoints $\mathcal{K}_{live}$ from the current observation $\mathcal{O}_{live}$, something not possible when using a single policy network. Going from RGB-D to a list of 3D points for inputs and outputs allows us to apply simple geometric data augmentation techniques for KAT, such as normalisation or random translations and rotations. This leads to a stronger spatial generalisation: in the "grasp a can" and "grasp a beer" tasks, we test performance of each method when the objects are on the table, or when they are positioned on top of other objects (a box, a microwave), a situation never encountered in the training videos. By running 5 test episodes for each case, we demonstrate how R+X retains strong performance, while the performance of the baselines drop. Results and an example of the predicted actions can be seen in Fig. 7.
Hard Language Generalisation: By leveraging the language and video understanding abilities of recent large Vision Language Models, such as Gemini, R+X can interpret and execute nuanced commands. To evaluate this, we set up a scene with many objects, and ask the robot to perform three tasks: "give me something to call my mom", "give me a non-alcoholic drink", and "make the room less dark". We run 5 test episodes for each of these commands, modifying the positions of the objects and the distractors. The language-conditioned policies struggle to interpret these commands, since they are strongly out of distribution. R+X, on the other hand, leverages Gemini’s ability to understand the meaning of these commands, as it is able to retrieve useful video clips from the human video (respectively, picking up the phone, grasping a Fanta can, and turning on the light). Results and an example of output gripper trajectories can be seen in Fig. 7.
Figure 8: Left (barplots 1-3): We compare R+X and the baselines’ ability to learn tasks in succession, and the time needed to learn such new tasks. Rightmost barplot: Gemini’s retrieval performance.
Can R+X learn task sequentially over time? In these experiments, we demonstrate how R+X can learn tasks continually, with no need for any additional training or finetuning, while obtaining strong performance both on the new task and on the old ones, a desirable ability for a robot learning from an ever increasing dataset of human experience. To measure this ability, and highlight the different behaviour with respect to a single language-conditioned policy, we first collect 10 demos for 3 tasks: grasping a can, cloth in basket, and picking up the phone. We train the baselines on these tasks (after extracting captions via Gemini as described for the experiments of Table 1) and evaluate them and R+X. Then, we add 10 demos of 3 new tasks (kettle on stove, open box, plate in rack), finetune the baselines, then measure performance on the 3 new tasks, and on the 3 old tasks, for all methods. In Fig. 8, Left, we see how the performance of R3M-DiffLang and Octo deteriorates on the old tasks, due to the well known effect of catastrophic forgetting [34]. On the other hand, R+X performance does not deteriorate, due to the ability to retrieve from an ever-growing video of tasks and adapting the behaviour model at test time on the retrieved data.
If instead we train the baselines on all the data each time, the performance does not deteriorate: however, this leads to a substantial growth in time needed to train. As the dataset grows, those networks would need an increasing amount of time per each new added task in order to be retrained, while R+X does not need any training or finetuning as further tasks are added to the video.
How does the length of the videos affect the retrieval performance? In this experiment, we first collect a video with 10 demos of 5 tasks (kettle on stove, cloth in basket, pick up phone, cloth in washing machine, close microwave), and measure Gemini’s ability to retrieve clips for each task. We then add 10 demos for 5 additional tasks, and compute again the precision and recall of retrieval of the original 5 tasks, to study how length of video and number of tasks affects it. The original video is close to 4 minutes in length, while the second is close to 10 minutes. In Fig. 8, Right, we show the mean over all tasks of the retrieval’s precision and recall. Results demonstrate how Gemini’s performance remains high while increasing the length of the video, allowing to scale R+X as the human records more and more tasks.
5 Discussion and Limitations
We have presented R+X, a framework to learn robot skills from long, unlabelled videos of humans interacting naturally in everyday environments, and demonstrated the clear benefit of using retrieval over training monolithic language-conditioned policy networks. There are, however, limitations which we now describe, opening up possible avenues for future work. The method is currently bottlenecked by the performance of Vision Language Model; however, such models have seen dramatic improvement in recent times and this will likely continue. While we can learn a large series of everyday tasks, the errors arising from the hand pose prediction currently inhibits us from learning precise tasks. Furthermore, while the keypoints extraction leads to strong spatial generalisation properties, it can become a bottleneck in the presence of many similar objects; methods that find a dynamic number of keypoints might tackle this issue. Finally, due to the use of Foundation Models for retrieval and execution, our method has a few seconds of delay between the issuing of the command and execution; however, this has reduced dramatically in recent months with the release of faster models such as GPT-4o [35] or Gemini Flash [24].