### v2

# R+X: Retrieval and Execution from Everyday Human Videos

Abstract
We present R+X, a framework which enables robots to learn skills from long, unlabelled, first-person videos of humans performing everyday tasks. Given a language command from a human, R+X first retrieves short video clips containing relevant behaviour, and then executes the skill by conditioning an in-context imitation learning method (KAT) on this behaviour. By leveraging a Vision Language Model (VLM) for retrieval, R+X does not require any manual annotation of the videos, and by leveraging in-context learning for execution, robots can perform commanded skills immediately, without requiring a period of training on the retrieved videos. Experiments studying a range of everyday household tasks show that R+X succeeds at translating unlabelled human videos into robust robot skills, and that R+X outperforms several recent alternative methods. Videos and code are available at https://www.robot-learning.uk/r-plus-x.
Given a language prompt, R+X first retrieves short relevant video clips extracted from a long unlabelled video of a human performing everyday tasks, recorded with a wearable camera. By using the retrieved video clips and a VLM, R+X performs in-context imitation learning, allowing it to immediately generate and execute the desired behaviour on the robot.
footnotetext: Joint first authorship. The Robot Learning Lab at Imperial College London. Contact at: {g.papagiannis21, n.di-palo20}@imperial.ac.uk
I Introduction
In this work, we study the problem of learning robot skills from long, unlabelled, first-person videos of humans performing everyday tasks in everyday environments. ”Long, unlabelled videos” means that a human simply goes by their everyday life and passively records videos of themselves without the need to specify which behaviour is being performed. ”Everyday environments” means that videos contain diversity in tasks, scenarios, objects, distractors, illumination, and camera viewpoints and motion. It is likely that such videos will become abundant through the adoption of wearable devices such as AR headsets and glasses [1, 2, 3, 4], and thus offer an unprecedented opportunity for future scalability if robots could learn from such videos, instead of learning exclusively from costly tele-operation.
Previous approaches to learning from videos of humans have often relied on a set of strong constraints and assumptions, such as human videos manually aligned with robot videos, human videos manually labelled with language descriptions or demonstrations on robot or MoCap hardware [5, 6, 7, 8, 9, 10, 11]. In this work, we remove all of these constraints completely and present a framework that requires only an unlabelled first-person video depicting tens of tasks.
Our framework, R+X, is a two-stage pipeline of Retrieval and EXecution shown in Figure R+X: Retrieval and Execution from Everyday Human Videos, which uses Foundation Models for both stages. Upon receiving a language command from a user, a Vision Language Model retrieves all clips where the human executes the specified task. By extracting the trajectories of the human’s hand in each clip, we then employ a few-shot in-context imitation learning method to condition on these trajectories, which enables a robot to ingest, learn from, and replicate the retrieved behaviours to previously unseen settings and objects. The recent literature demonstrated that, by finetuning large Vision Language Models on robotics data, they can transfer their common knowledge and ability to reason and plan to robotics settings [12, 13]. This, however, requires very expensive finetuning of often intractably large models. With our proposed framework, we can equally leverage these abilities but now via retrieval and video understanding, thus without the need for any finetuning.
In summary, the two main properties of R+X are: 1) it enables robots to execute everyday tasks from language commands, given long, unlabelled, first-person videos collected naturally by recording a human’s daily activities, and 2) it achieves this without the need for any training or finetuning of models, allowing it to learn and execute tasks immediately, while still inhereting many strengths of large Vision Language Models.
II Related Work
Figure 1: The main assumptions and constraints of many recent Learning from Observation methods.
Collecting Robotics Data. Scaling data collection has been proven to be a successful path towards increasingly more general machine learning models [14, 15, 16, 17]. To collect robotics data, the most common paradigm is to teleoperate robots, collecting datasets of paired observations and actions [18]. This, however, needs dedicated teleoperation hardware, and needs human operators to allocate their time to actively teach a robot new tasks. In our framework, a human user interacts with their environments as usual, completing the tasks they wish, while a robot passively learns to emulate it, resulting in a more scalable and time efficient paradigm.
Learning from Human Videos. Many recent works have proposed solutions to teach new skills to robots by observing a human executing such tasks. However, as we illustrate in Fig. 1, they often relied on a set of assumptions and constraints that limited their use ”in-the-wild”. [6, 19] require a combination of human data and either robot exploration or teleoperation, therefore needing active assistance of the user in teaching the robot, and videos are recorded from a fixed, third-person camera. R+X relies entirely on the videos recorded by the user with a mobile camera in their natural environments. [8, 10, 9] can learn robot skills entirely from human videos. [8, 10], however, focus on replicating the object trajectory from the demo, and cannot perform tasks that do not involve grasping and moving, such as pushing or pressing, that R+X can execute. [9] can learn a larger repertoire of skills, but still relies on learning single tasks in isolation from a fixed camera. Instead, R+X does not assume a fixed camera and can replicate tasks from a given language command after receiving a single, long, unlabelled video depicting tens of tasks, without the need to specify which clip demonstrated which behaviour. Methods like [11] allow a user to naturally interact with their environment while collecting data that can teach a robot new skills. However, they require additional hardware to wear, like specialised MoCap gloves. R+X only needs a single RGB-D camera, and does not require any extrinsics calibration. This means that data could be recorded using a wearable camera, smart glasses, AR visors, and more.
Language-Conditioned Policy Learning. Language is considered a viable way to instruct robots and guide their task execution [18, 13, 12]. The common approach is to train a large, language and image conditioned policy that can, at test time, receive user commands and camera observations [18]. This, however, presents some criticalities: training such models can require enormous computational effort [12], and unlike other machine learning applications, robots should learn new skills over time. Therefore, re-training or finetuning such networks should be avoided. R+X therefore does not train or finetune any networks, but takes advantage of pre-trained models able to retrieve examples of tasks [15], and to learn to emulate and execute new behaviour directly at deployment via in-context learning [20].
Additional papers are discussed in the appendix.
Figure 2: Upon receiving a language command, R+X retrieves all the relevant clips from the human video. Each retrieved clip is transformed from pixels to a sparse 3D points representations of the hand joints movement and salient parts of the visual observation.
III R+X: Retrieval and Execution
We now describe R+X, our proposed method to learn robot skills from long, unlabelled videos of humans interacting with their environments. Additional details accompanying our method can be found in the appendix and on our website. We assume R+X has access to a long, unlabelled video of a person performing a multitude of tasks in many different locations. We call this long, first-person video of everyday tasks the ”human $\mathcal{H}$ in the rest of the paper. The goal of our method is to learn to emulate the behaviours recorded in the human video upon receiving a language command.
Overall, from a high-level perspective, our pipeline takes three inputs: 1) a single, long, unlabelled video of all the interactions recorded by the humans, $\mathcal{H}$ 2) a task to execute in language form, $\mathcal{L}$ and 3) the current observation of the robot as an RGB-D image, $\mathcal{O}_{live}$. It then outputs a trajectory of 6-DoF gripper poses, which are executed by the robot to tackle the task at hand.
There are however a set of non-trivial challenges: while receiving a language command to execute at deployment, the robot receives no language information before deployment: the recorded video contains no more information than the recorded visual frames. Additionally, the robot receives no action information: unlike the case of teleoperation, no joints or end-effector position/velocity/acceleration data are recorded, and the correct movements need to be inferred by the videos alone: the problem is additionally complex due to the cross-embodiment between human videos and final robot actions. Furthermore, as the user is interacting with their natural environment, the visual observations can be filled with distractor objects typical of household and offices, unrelated to the task the user is performing. To tackle these challenges, we leverage the abilities of Foundation Models.
Specifically, at the first phase of R+X, which we call the retrieval phase, we use a VLM to extract from the human video all the examples of the desired behaviour described in the prompt $\mathcal{L}$ as a list of $Z$ shorter video clips $[\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{Z}]$. We map the $Z$ videos into a lower dimensional 3D representation, extracting for each video: (1) a list of $K$ visual 3D keypoints that describe the scene, $[k_{1},k_{2},\dots,k_{K}]$ where $k=(x_{k},y_{k},z_{k})$, and (2) the movement of the user’s hand as a trajectory of length $T$ of 21 hand joints that parametrise the MANO hand model [21], $\mathcal{J}=[j_{1},j_{2},\dots,j_{T}]$, where $j=[[x_{0},y_{0},z_{0}],\dots,[x_{21},y_{21},z_{21}]]$. Finally, at the second stage of R+X, which we call the execution phase, to emulate the behaviours observed in the retrieved video clips, we condition a few-shot in-context imitation learning model on this data that, given a live observation of the environment, it generates a trajectory of 3D hand joints to execute the desired behaviour described in the prompt. To map such joints to gripper poses, we develop a custom procedure that we describe in the appendix.
Figure 3: The visual 3D keypoints of the first frame of each of the Z videos obtained from the retrieval phase along with each extracted hand joint trajectory are used as context for KAT. To execute a skill, visual 3D keypoints are extracted from the live observation and used as input to KAT which generates a sequence of hand joints. By mapping this sequence to gripper poses the robot executes the desired task.
III-A Retrieval: Extracting visual examples from a long, unlabelled video
The first main phase of R+X is retrieval, shown in Figure 2. Upon receiving a language command $\mathcal{L}$, and given the human video $\mathcal{H}$, the goal is to retrieve all video clips from the human video that depict the execution of the requested task. This is accomplished using a recent Vision Language Model (VLM), Gemini Pro 1.5 Flash [22]. Gemini, which we denote $\mathcal{G}$, is natively multi-modal and can take as input images, videos, and text, and outputs text. We prompt the model with the human video, and ask it to retrieve the starting and ending seconds at which the received task happens. The inputs of this phase is therefore a language command and the human video, and the output is a list of $Z$ shorter video clips demonstrating the desired task, $\mathcal{G}(\mathcal{H},\mathcal{L})\rightarrow[\mathcal{V}_{1},\dots,\mathcal{V}_{Z}]$. Each clip comprises $T$ RGB-D frames, where $T$ can vary across different clips.
III-A1 Preprocessing Videos into a Sparse 3D Representation
Given the $Z$ extracted video clips, $[\mathcal{V}_{1},\dots,\mathcal{V}_{Z}]$, we apply a preprocessing step that converts each video clip from a list of RGB-D frames to a set of 3D points describing the visual scene and the trajectory of the human’s hand, a representation that we will then feed to our few-shot in-context imitation learning model.
Visual 3D keypoints. To transform the complex RGB-D observations into a lower-dimensional, easier to interpret input, we harness the powerful vision representations generated by DINO [23], a recent Vision Foundation Model. As proposed in [24, 20], given the $Z$ clips retrieved as described before, we find a set of $K$ common 3D visual keypoints that capture interesting semantic or geometrical aspects of the scene. We extract these keypoints from the first frame of each of the $Z$ videos only.
We first compute the DINO descriptors for the first frame of each of the $Z$ videos, obtaining a list of $Z$ different $N\times 384$ outputs, where $N$ is the number of patches of each image [25, 26, 23]. Via a clustering and matching algorithm [24], we select from the $N\times 384$ descriptors of the first frame of the first video, $\mathcal{O}_{\mathcal{V}_{1},1}$, a list of the $K$ descriptors that are the most common in all the remaining $Z-1$ frames. We denote these descriptors as $\mathcal{D}\in\mathbb{R}^{K\times 384}$. Therefore, this way, we autonomously extract descriptors that focus on the object of interest, as its appearance is common among videos, while distractors and overall scene will vary [24, 20].
Finally, for each of the $K$ descriptors in $\mathcal{D}$, we extract keypoints by finding the $K$ nearest neighbours between the $N\times 384$ descriptors of each the $Z$ frames, and compute their 2D coordinates. We then project these in 3D using each frame’s depth image and known camera intrinsics. As such, given the retrieved clips, we obtain and store a list $\Lambda=[\mathcal{K}_{V_{1}},\dots,\mathcal{K}_{V_{z}}]\in\mathbb{R}^{Z\times K\times 3}$ of visual 3D keypoints, $K$ for each clip, where each $\mathcal{K}_{V_{i}}=[k_{1},...,k_{K}]$ and $k_{j}=(x_{j},y_{j},z_{j})$. For a more detailed description please refer to [24].
Hand Joint Actions. To extract human actions from each retrieved video clip we use the HaMeR model [27], a recent technique for 3D hand pose estimation from images based on DINO. In particular, using HaMeR we extract from each video frame $\mathcal{O}_{\mathcal{V}_{z},t}$ (where $1\leq t\leq T$) of each of the $Z$ clips the 3D hand pose, represented as a set of 21 3D points, $j_{\mathcal{V}_{z},t}$, describing the hand joints that parameterise the MANO hand model, as it is commonly done in the literature [21]. As the camera moves between frames, we design a stabilisation technique to compute transformations between camera poses that is robust to dynamic scenes (for more details please see the appendix). This enables us to express the extracted hand joints relative to a single reference frame, that of the first camera frame of each clip. For each video clip $\mathcal{V}_{z}$, this process results in a sequence of 3D hand joint actions $\mathcal{J}_{\mathcal{V}_{z}}=[j_{\mathcal{V}_{z},1},\dots,j_{\mathcal{V}_{z},T}]$, expressed relative to the first frame of each video $\mathcal{O}_{\mathcal{V}_{z},1}$ where the visual 3D keypoints are also expressed in. As such, at the end of this process we are left with a list $\mathcal{M}=[\mathcal{J}_{\mathcal{V}_{1}},...\mathcal{J}_{\mathcal{V}_{Z}}]$ of $Z$ hand joint action sequences.
In summary, from the $Z$ retrieved video clips, we extracted the list of hand joints actions $\mathcal{M}$, along with the list of visual 3D keypoints $\Lambda$, that will be used as context for our in-context imitation learning model, as $Z$ input-output pairs.
III-B Execution: Few-Shot, In-Context Imitation from Video Examples
The second main phase of R+X is execution. Our framework is based on the use of a model capable of performing few-shot, in-context imitation learning, receiving a few examples of desired inputs and outputs pairs describing a desired behaviour, and able to replicate such behaviour immediately upon receiving a new input. We use Keypoint Action Tokens (KAT) [20] to achieve this, a recently proposed technique that takes 3D visual keypoints as input and outputs a trajectory of 3D points describing the gripper movement. Instead of explicitly training a model on robot data, KAT demonstrates that recent, off-the-shelf Large Language Models are able to extract such numerical patterns, and behave as few-shot, in-context imitation learning machines, without the need for any further finetuning.
Figure 4: We test R+X on 12 everyday tasks, executed by a human in different rooms and with different distractors.
Given the output of the retrieval phase $[\Lambda,\mathcal{M}]$ and a new, live RGB-D observation collected by the robot $\mathcal{O}_{live}$, we extract its visual 3D keypoints representation $\mathcal{K}_{live}$ by first extracting its $N\times 384$ DINO descriptors, and then finding the $K$ nearest neighbours to each of the $K$ descriptors in $\mathcal{D}$, that we obtained in the retrieval phase. This results into $K$ 2D coordinates, that we project in 3D. We then input to KAT as context $[\Lambda,\mathcal{M}]$, as $Z$ examples of the desired input-output mapping, and the new visual 3D keypoints, and generate a new trajectory of desired hand joint actions, $\text{KAT}([\Lambda,\mathcal{M}],\mathcal{K}_{live})\rightarrow\mathcal{J}_{live}$ with $\mathcal{J}_{live}=[j_{1},\dots,j_{T}]$, as shown in Figure 3. We then map the predicted trajectories of hand joints actions into gripper poses as described in our website and appendix, and execute them.
To summarize, given a language command $\mathcal{L}$ and a human video $\mathcal{H}$, R+X first retrieves $Z$ videos depicting the described behaviour using a VLM. From the retrieved videos, a list of $[\Lambda,\mathcal{M}]$ visual 3D points and hand joint actions are extracted. Then, to execute the desired behaviour described in the prompt $\mathcal{L}$, $[\Lambda,\mathcal{M}]$ along with the visual 3D keypoints of the live observation $\mathcal{K}_{live}$ are used as context for KAT that performs few-shot in-context imitation learning to generate a trajectory of hand joints actions, which is mapped to gripper poses and executed on the robot.
IV Experiments
Human Video. We collect the human video $\mathcal{H}$ using an Intel RealSense 455, worn by a human on their chest as shown in Figure R+X: Retrieval and Execution from Everyday Human Videos. To reduce downstream computational time, we filter out each frame in which human hands are not visible right after recording. As our robot is single-armed, we limit ourselves to single hand tasks. However, our method could identically be applied to bimanual settings and dexterous manipulators. The video is collected in many different rooms and buildings.
Robot Setup. At execution, we use a Sawyer robot equipped with a RealSense 415 head-camera. The robot is equipped with a two-fingered parallel gripper, the Robotiq 2F-85. As the robot is not mobile, we setup different scenes in front of it with variations of the tasks recorded by the human, placing several different distractors for each task, while the human video was recorded in many different rooms. Although we have a wrist-camera mounted, we do not use that in our work.
Tasks. To evaluate our proposed framework, we use a set of 12 everyday tasks, where a human interacts with a series of common household objects, listed in Fig. 4. We include movements like grasping, opening, inserting, pushing, pressing, and wiping.
Baselines. We compare R+X, and its retrieval and execution design, to training a single, language-conditioned policy. To obtain language captions from the human video, we use Gemini to autonomously caption snippets of the video, obtaining a (observation, actions, language) dataset. We finetune R3M (ResNet-50 version [28]) [29] and Octo [30] on this data. We extend R3M to also encode language via SentenceBERT and use a Diffusion Policy [31] head to predict actions from intermediate representations. We denote this version as R3M-DiffLang.
Videos, code, and additional implementation details can be found on our website: https://www.robot-learning.uk/r-plus-x
Method / Task Plate Push Wipe Beer Wash Box Kettle Micro. Basket Phone Can Light Avg.
R3M-DiffLang 0.5 0.7 0.4 0.7 0.5 0.5 0.4 0.8 0.7 0.4 0.7 0.3 0.55
Octo 0.5 0.8 0.5 0.6 0.5 0.5 0.4 0.7 0.6 0.4 0.6 0.3 0.53
R+X 0.6 0.8 0.7 0.8 0.6 0.7 0.6 0.8 0.7 0.7 0.8 0.6 0.7
TABLE I: Result of the various methods on the 12 proposed tasks.
IV-A Results
Can R+X learn robot skills from long, unlabelled videos? How does it perform with respect to a monolithic language-conditioned policy? In these experiments, we evaluate the performance of R+X in learning a set of everyday skills. At deployment, we place the object to be interacted with in front of the robot, and issue a language command. We then run the retrieval and execution pipeline using the human video, the current observation and the language command. We run 10 evaluations per task, randomising 1) the object pose 2) the type and number of distractors 3) for tasks where it is possible, we swap the object to interact with, with another one from the same class to test for generalisation (e.g. a different can, a different piece of clothing, a different telephone). We provide more details in the appendix.
In Table I, we report the performance of R+X on these tasks, together with the baselines. As we demonstrate, the framework is able to tackle a wide range of everyday tasks, surpassing the monolithic policy baselines. These results prove the benefit of modeling language-conditioned learning from observation as distinct retrieval and execution phases, to fully leverage the abilities of recent Foundation Models. In the appendix, we investigate more in depth the performance of R+X on unseen objects and in the presence of distractors.
Figure 5: Examples of spatial, language and distractors generalisation. Gripper trajectories move from red to blue.
What are the main sources of difference in performance between R+X and a monolithic policy? In these experiments, we investigate what changes in the inputs lead to the most noticeable difference in performance between R+X and the baselines, R3M-DiffLang and Octo. We explored two aspects, related to two properties of R+X:
Hard Spatial Generalisation: R+X, by retrieving videos of the desired task, can also extract a series of relevant keypoints $\mathcal{K}_{live}$ from the current observation $\mathcal{O}_{live}$, something not possible when using a single policy network. Going from RGB-D to a list of 3D points for inputs and outputs allows us to apply simple geometric data augmentation techniques for KAT, such as normalisation or random translations and rotations. This leads to stronger spatial generalisation: in the ”grasp a can” and ”grasp a beer” tasks we test each method’s performance when the objects are on the table, or when they are positioned on top of other objects (a box, a microwave). By running 5 test rollouts for each case, we demonstrate how R+X retains strong performance, while the baselines’ performance drops. Results and an example of the predicted gripper trajectories can be seen in Fig. 5, middle.
Hard Language Generalisation: By leveraging the language and video understanding abilities of recent large Vision Language Models, such as Gemini, R+X can interpret and execute nuanced commands. To evaluate this, we setup a scene with many objects, and ask the robot to perform three tasks: ”give me something to call my mom”, ”give me a non-alcoholic drink” and ”make the room less dark”. We run 5 test rollouts for each of these commands, modifying the position of the objects and the distractors. The language-conditioned policies struggle to interpret these commands, that are strongly out of distribution. R+X, on the other hand, leverages Gemini’s ability to understand the meaning of these commands, as it is able to retrieve useful video clips from the human video (respectively, picking up the phone, grasping a Fanta can, and turning on the light). Results and an example of output gripper trajectories can be seen in Fig. 5, left.
Figure 6: We compare R+X and the baselines’ ability to learn tasks in succession, and the time needed to learn such new tasks.
Distractors Generalisation: One of the main differences between R+X and the baselines is that the former, after retrieving a set of video clips from $\mathcal{H}$ depicting the requested task, can extract a set of keypoints that are semantically and geometrically meaningful for the objects to interact with, as described in the main paper. These keypoints are generally robust to distractors, as they focus only on the DINO features that are common in all images, whereas distractors typically vary throughout the human video $\mathcal{H}$ and the execution. The baselines, being monolithic policies, receive as input the observation as it is, and must learn to generalise to the presence of distractors during the long training phase.
To evaluate the effect of distractors on the scene, we test R+X and the baselines on two tasks, ”grasp a can” and ”put the cloth in the basket”, emulating the experimental scenario of the ”Hard Generalisations” sections. We run 10 runs without distractors, and 10 with, explicitly measuring the performance in the two cases. We can see in the results of Fig. 5 right, how R+X is more robust to the presence of random distractor objects, highlighting the advantage of extracting semantically and geometrically meaningful keypoints after having received a language command and having retrieved the corresponding video clips.
Can R+X learn task sequentially over time? In these experiments, we demonstrate how R+X can learn tasks continually, with no need for any additional training or finetuning, while obtaining strong performance both on the new task and on the old ones, a desirable ability for a robot learning from an ever increasing dataset of human experience. To measure this ability, and highlight the difference behaviour with respect to a single language-conditioned policy, we first collect 10 demos for 3 tasks. We train the baselines on these tasks (after extracting captions via Gemini as described for the experiments of Table I) and evaluate them and R+X. Then, we add 10 demos of 3 new tasks, finetune the baselines, then measure performance on the 3 new tasks, and on the 3 old tasks for all methods. In Fig. 6, we see how the performance of R3M-DiffLang and Octo deteriorates on the old task, due to the well known effect of catastrophic forgetting [32]. On the other hand, R+X performance does not deteriorate, due to the ability to retrieve from an ever-growing video of tasks and adapting the behaviour model at test time on the retrieved data.
If instead we train the baselines on all the data each time, the performance does not deteriorate: however, this leads to a substantial growth in time needed to train. As the dataset grows, those networks would need an increasing amount of time per each new added task in order to be retrained, while R+X does not need any training or finetuning.
V Conclusion
We presented R+X, a method to learn robot skills from long, unlabelled videos of humans interacting with their environments. Given a language command and a live observation of the scene, R+X retrieves a sequence of relevant video clips from a long unlabelled video which it uses as context for an in-context imitation learning method that predicts what actions to execute. Through several real-world experiments on 12 everyday tasks, we demonstrate that R+X can successfully generalise to previously unseen objects, scenes, and object placements, is robust to distractor objects, and its use of VLMs for video retrieval enables R+X to generalize to complex language commands demonstrating its clear benefits over training monolithic language-conditioned policy networks. For a discussion on limitations and future work please see the appendix.